# Optimizing entanglement in two-qubit systems

### v1

@@ 2.2 Entanglement conditions
The transposition of the X–state (5) corresponds to make $r_{14}\rightarrow r_{23}$. Eigenvalues (A-1) and eigenvectors (A-2) must be transformed accordingly to solve the eigenvalue equation $\rho_{X}^{T_{2}}|\widetilde{\epsilon}\,\rangle=\widetilde{\Lambda}|\widetilde{\epsilon}\,\rangle$. Then, it is found that $\rho_{X}^{T_{2}}$ is positive semidefinite whenever $|r_{23}|\leq\sqrt{r_{11}r_{44}}$ and $|r_{14}|\leq\sqrt{r_{22}r_{33}}$. Applying the Peres–Horodecki criterion [14, 15], we find that $\rho_{X}$ is entangled as long as one of the following inequalities hold
@@ 3 Convex optimization
Consider for example the Werner–states (10). If $k=1,2$, we find $y=0$, $x=\frac{1}{2}|q|$, $y_{0}=\frac{1}{4}(1-q)$ and $x_{0}=\frac{1}{4}(1+q)$, see Figure 2(a). To find entanglement, $x$ must obey inequalities (15). If $k=3,4$, the results are obtained from the latter by changing $x_{0}\leftrightarrow y_{0}$ and $x\leftrightarrow y$, see Figure 2(b). Entanglement is now found whenever $y$ obeys inequalities (16). The most striking thing about these results is that all states $\rho_{W_{k}}$ share the same entanglement profile, they are entangled for $q\in(1/3,1]$, regardless of $k$.
@@ 3.1 Entanglement
$\bullet$ Region $\mathcal{M}_{0}$. No entanglement is allowed if $r_{11}r_{44}=r_{22}r_{33}\neq 0$. That is, none of inequalities (15)-(16) apply if $x_{0}=y_{0}\neq 0$. Instead we have $0\leq x,y\leq x_{0}$, so we have a square $\square_{x_{0}}$ of sides $x_{0}$, with $x_{0\max}=1/4$, see Figure 4. These points produce separable mixed states $\rho_{X}$ that can be identified by their rank. According to the eigenvalues (A-1), the vertex $q_{\star}=(x_{0},x_{0})$ represents rank-2 states. In turn, $(x_{0},y)$ and $(x,y_{0})$, located respectively along the right-vertical and top sides of the square, define rank-3 states. Any other point of $\square_{x_{0}}$ gives rank-4 states.
@@
In general, by extending the domain of $x$ to the maximum we optimize the entanglement in $\mathcal{M}_{x}$. For this, it is enough to set $y_{0}=0$. We find three different states
@@
The classification developed above is very close to the parametrization of two-qubit X–density matrices with a fixed rank reported in Section 2.1 of [26]. The latter, based on the coefficients that define the characteristic equation for the density operator, considers the Newton-Girard formulae [34] and a very concrete parametrization of the matrix-elements of $\rho_{X}$. Beyond using the computational basis, our treatment is free of parametrization for the matrix elements $r_{jk}$, so it aims to be more practical and, therefore, universal. Nevertheless, the parameters used in [26] are very valuable as they comprise the information needed to obtain a unitary transformation that converts arbitrary two-qubit states into their X counterparts.
@@ 3.2 Geometric $L$-measure
Let us apply the measure $L$ to the points $p\in\mathcal{S}$ that define the states classified in the previous section. States $\rho_{\mathcal{M}_{1}}$ are such that $\inf d(q,p)$ is given by the points $q=(y_{0},y)$, see Figure 7(a). Therefore
@@
If $x$ and $y$ take the maximum value in their respective domains we have
@@
The most notable of the previous results is that the mathematical expression of the geometric measure $L$ coincides precisely with the Hill-Wootters concurrence calculated for the $\rho_{X}$ states. In fact, according with [28, 29], the concurrence for any state $\rho$ is given by $C(\rho)=\max\{0,\,\sqrt{\ell_{1}}-\sqrt{\ell_{2}}-\sqrt{\ell_{3}}-\sqrt{\ell_{4}}\}$. The $\ell_{k}$’s are the eigenvalues (in decreasing order) of the matrix $R_{\rho}=\rho\left(\sigma_{2}\otimes\sigma_{2}\right)\rho^{*}\left(\sigma_{2}\otimes\sigma_{2}\right)$, with $\sigma_{2}$ the well-known anti-diagonal Pauli-matrix.
@@
which once summarized leads to the identity $C(\rho_{X})=L(\rho_{X})$.
@@ 3.3 Determination of entanglement
The entanglement of the Werner–states $\rho_{W_{k}}$, quantified by $L$ and $\varepsilon$, is shown in Figure 8(a) (see also Figure 2). The optimization is achieved in half of the domain of $q$ only.
@@
Figure 8: The geometric measure $L$ (continuous curve) and entanglement of formation $\varepsilon$ (dotted curve) applied to (a) the Werner–states $\rho_{W_{k}}$ and (b) the statistical mixture $\rho_{B}$ of only two Bell–states. There is no entanglement for half the domain of $q$ in $\rho_{W_{k}}$, compare with Figure 2. In turn, states $\rho_{B}$ lack entanglement only at one point ($q=1/2$) of the corresponding domain, compare with Figures 3(a) and 3(d).
Figure 8(b) shows the result of applying $L$ and $\varepsilon$ to the convex combinations of Bell–states exhibited in Figures 3(a) and 3(d). This time the optimization is over almost the entire domain of $q$, mainly in the neighborhoods of $q=0$ and $q=1$, which serve as accumulation points. These systems are separable only for $q=1/2$.
@@
Figure 9: The geometric measure $L$ (continuous curve) and entanglement of formation $\varepsilon$ (dotted curve) applied to the generalized Werner–states shown in Figure 3. The case depicted in Figure 8(b) is recovered if $s=1$. In general, the maximum values of $L$ and $\varepsilon$ decrease with $s$ while the region where they become zero increases.
@@
Figure 10: (a) The geometric measure $L$ applied to the time-dependent diatomic state $\rho_{\operatorname{2at}}$. As time passes, $L$ decreases from 1 to zero and then increases until it is almost equal to 1. This process is repeated over and over again (b) Comparison of $L(\rho_{\operatorname{2at}})$ with the difference $y-x_{0}$. Clearly, inequality $y>x_{0}$ is violated in the same time intervals where $L=0$.
@@
The envelope formed by the minima of S establishes an upper limit for the entanglement measurement. To our knowledge, this surprising property of entropy shared by reduced states has not been observed in the literature on the matter.
@@
Figure 11: The envelope formed by the minima of the one-qubit entropy (blue curve) defines an upper bound for entanglement in the time-dependent diatomic state $\rho_{\operatorname{2at}}$ (a) The geometric measure $L$ is red-shaded (b) The entanglement of formation $\epsilon$ is purple-shaded.
@@ 4 Conclusions
With the above results we have constructed a convex representation that facilitates the study of entanglement in two-qubit systems. The description is purely geometric and is based on two of the most striking results we have found:
@@
The $L$-measure coincides with the Hill–Wootters generalized concurrence $C$ for states $\rho_{X}$. That is, our results give $C$ a geometric meaning that is not obvious without the convex structure introduced in this work. What is even better is that, unlike $C$, the geometric measure $L$ distinguishes the rank of states with maximum entanglement.
@@
The applicability of our method goes far beyond the time-independent systems typically discussed in the literature. We have also studied what happens to entanglement when the system depends explicitly on time. Within the geometric representation, the time-evolution of $\rho_{X}$ describes paths that transit between areas with different entanglement strength, even invading regions where the state becomes separable to return to the entanglement zones, and so on. The latter opens the possibility of manipulating and controlling entanglement in two-qubit systems by using inverse problem techniques [39, 13].
We have found another outstanding result for the time-dependent system studied here. The envelope formed by the minima of $S$, the entropy of the reduced states, establishes an upper bound for the entanglement measurement. To our knowledge, this surprising property has not been reported in the literature on the matter.

### v2

@@ 2.2 Entanglement conditions
The transposition of the X–state (5) corresponds to make $r_{14}\rightarrow r_{23}$. Eigenvalues (A-1) and eigenvectors (A-2) must be transformed accordingly to solve the eigenvalue equation $\rho_{X}^{T_{2}}|\widetilde{\epsilon}\,\rangle=\widetilde{\Lambda}|\widetilde{\epsilon}\,\rangle$. Then, it is found that $\rho_{X}^{T_{2}}$ is positive semidefinite whenever $|r_{23}|\leq\sqrt{r_{11}r_{44}}$ and $|r_{14}|\leq\sqrt{r_{22}r_{33}}$. Applying the Peres–Horodecki criterion [19, 20], we find that $\rho_{X}$ is entangled as long as one of the following inequalities hold [37]:
@@ 3 Convex optimization
Consider for example the Werner–states (13). If $k=1,2$, we find $y=0$, $x=\frac{1}{2}|q|$, $y_{0}=\frac{1}{4}(1-q)$ and $x_{0}=\frac{1}{4}(1+q)$, see Figure 2(a). To find entanglement, $x$ must obey inequalities (28). If $k=3,4$, the results are obtained from the latter by changing $x_{0}\leftrightarrow y_{0}$ and $x\leftrightarrow y$, see Figure 2(b). Entanglement is now found whenever $y$ obeys inequalities (29).
@@
The most striking thing about the above results is that all states $\rho_{W_{k}}$ share the same entanglement profile: they are entangled for $q\in(1/3,1]$, regardless of $k$.
@@ 3.1 Entanglement
$\bullet$ Region $\mathcal{M}_{0}$. No entanglement is allowed if $r_{11}r_{44}=r_{22}r_{33}\neq 0$. That is, none of inequalities (28)-(29) apply if $x_{0}=y_{0}\neq 0$. Instead we have $0\leq x,y\leq x_{0}$, so we arrive at a square $\square_{x_{0}}$ of sides $x_{0}$, with $x_{0\max}=1/4$, see Figure 4. These points produce separable mixed states $\rho_{X}$ that can be identified by their rank. According to the eigenvalues (A-1), the vertex $q_{\star}=(x_{0},x_{0})$ represents rank-2 states. In turn, $(x_{0},y)$ and $(x,y_{0})$, located respectively along the right-vertical and top sides of the square, define rank-3 states. Any other point of $\square_{x_{0}}$ gives rank-4 states.
@@
In general, by extending the domain of $x$ to the maximum we optimize the entanglement in $\mathcal{M}_{x}$. For this, it is enough to set $y_{0}=0$. We find three different sets of states
@@
The classification developed above is very close to the parametrization of two-qubit X–density matrices with a fixed rank reported in Section 2.1 of [38]. The latter, based on the coefficients that define the characteristic equation for the density operator, considers the Newton-Girard formulae [52] and a very concrete parametrization of the matrix-elements of $\rho_{X}$. Beyond using the computational basis, our treatment is free of parametrization for the matrix elements $r_{jk}$, so it aims to be more practical and, therefore, universal. Nevertheless, the parameters used in [38] are very valuable as they comprise the information needed to search for a unitary transformation that converts arbitrary two-qubit states into their X counterparts.
@@ 3.2 Geometric $L$-measure
Let us apply the measure $L$ to the points $p\in\mathcal{S}$ that define the states classified in the previous section.
States $\rho_{\mathcal{M}_{1}}$ are such that $\inf d(q,p)$ is given by the points $q=(y_{0},y)$, see Figure 7(a). Therefore
@@
As for systems composed of subsystems with equal entropy, see Section 2.1.1, we have immediate results:
$L(\rho_{B})=2\max\left\{0,\frac{|b_{1}-b_{2}|}{2}-\frac{(b_{3}+b_{4})}{2},\frac{|b_{3}-b_{4}|}{2}-\frac{(b_{1}+b_{2})}{2}\right\},$ (35)
$L(\rho_{W_{k}})=2\max\left\{0,\frac{|q|}{2}-\frac{|1-q|}{4}\right\},$ (36)
$L(\rho_{S})=2\max\left\{0,\frac{|q_{1}-q_{2}|}{2}-\varrho_{22},\frac{|q_{3}-q_{4}|}{2}-\varrho_{11}\right\}.$ (37)
To illustrate the applicability of the above expressions let us consider the state (11)-(12). From (35) one has
$\left.L(\rho_{B})\right|_{b_{3}=b_{4}=\kappa}=2\max\left\{0,\frac{|b_{1}-b_{2}|}{2}-\kappa\right\}.$ (38)
Therefore, $\left.\rho_{B}\right|_{b_{3}=b_{4}=\kappa}$ is disentangled if $|b_{1}-b_{2}|\leq 2\kappa=1-(b_{1}+b_{2})$. For these parameters the interference between $|e_{1}\rangle=|00\rangle$ and $|e_{4}\rangle=|11\rangle$ is so small that the system ends up behaving like a statistical mixture (a classical state) of the base states. On the other hand, for $|b_{1}-b_{2}|>2\kappa$ we find entanglement with measure $\left.L(\rho_{B})\right|_{b_{3}=b_{4}=\kappa}=|b_{1}-b_{2}|-2\kappa$ The extreme condition $\kappa=0$ shows that the statistical mixture of the Bell–states $|\beta_{1}\rangle$ and $|\beta_{2}\rangle$ is entangled, with a measure $L$ that is greater as the interference between $|00\rangle$ and $|11\rangle$ is more intense. In particular, setting $b_{1}=q$ and $b_{2}=1-q$ we have $\left.L(\rho_{B})\right|_{b_{3}=b_{4}=0}=|2q-1|$. So state $\left.\rho_{B}\right|_{b_{3}=b_{4}=0}$ is disentangled for $q=1/2$ ($b_{1}=b_{2}$), as it is shown in Figures 3(a) and 3(d).
In general, if $x$ and $y$ take the maximum value in their respective domains we have
@@
The most notable of the previous results is that the mathematical expression of the geometric measure $L$ coincides precisely with the Hill-Wootters concurrence calculated for the $\rho_{X}$ states [35, 36, 37]. In fact, according with [45, 46], the concurrence for any state $\rho$ is given by $C(\rho)=\max\{0,\,\sqrt{\ell_{1}}-\sqrt{\ell_{2}}-\sqrt{\ell_{3}}-\sqrt{\ell_{4}}\}$. The $\ell_{k}$’s are the eigenvalues (in decreasing order) of the matrix $R_{\rho}=\rho\left(\sigma_{2}\otimes\sigma_{2}\right)\rho^{*}\left(\sigma_{2}\otimes\sigma_{2}\right)$, with $\sigma_{2}$ the well-known anti-diagonal Pauli-matrix.
@@
which once summarized is written as follows [35, 36, 37]:
$C(\rho_{X})=2\max\left\{0,|r_{14}|-\sqrt{r_{22}r_{33}},\,|r_{23}|-\sqrt{r_{11}r_{44}}\,\right\}.$ (39)
Comparing (39) with (34) we find $C(\rho_{X})=L(\rho_{X})$.
@@
In this context, the $L$-measure of Werner–states (36) deserves special attention. First of all, this verifies that all states $\rho_{W_{k}}$ share the same entanglement profile: they are entangled for $q\in(1/3,1]$, regardless of $k$, see Figure 2. On the other hand, two-qubit Werner–states can be prepared experimentally in such a way that the related hierarchy of quantum correlations can also be demonstrated. Indeed, in [27, 29], states (13) are prepared for various values of the mixing parameter $q$, so that concurrence $C(\rho_{W_{k}})=L({W_{k}})$ is measured with good agreement of (36).
@@ 3.3 Determination of entanglement
The entanglement of the Werner–states $\rho_{W_{k}}$, quantified by $L$ and $\varepsilon$, is shown in Figure 8(a) (see also Figure 2). The optimization is achieved in half of the domain of $q$ only. The experimental realization of Werner–states and the measurement of their entanglement can be found in [27, 29].
@@
Figure 8: The geometric measure $L$ (dotted curve) and entanglement of formation $\varepsilon$ (continuous curve) applied to (a) the Werner–states $\rho_{W_{k}}$ and (b) the statistical mixture $\rho_{B}$ of only two Bell–states. There is no entanglement for half the domain of $q$ in $\rho_{W_{k}}$, compare with Figure 2. In turn, states $\rho_{B}$ lack entanglement only at one point ($q=1/2$) of the corresponding domain, compare with Figures 3(a) and 3(d).
Figure 8(b) shows the result of applying $L$ and $\varepsilon$ to the convex combinations $\rho_{B}$ of Bell–states exhibited in Figures 3(a) and 3(d). This time the optimization is over almost the entire domain of $q$, mainly in the neighborhoods of $q=0$ and $q=1$, which serve as accumulation points. These systems are separable only for $q=1/2$.
@@
Figure 9: The geometric measure $L$ (dotted curve) and entanglement of formation $\varepsilon$ (continuous curve) applied to the generalized Werner–states shown in Figure 3. The case depicted in Figure 8(b) is recovered if $s=1$. In general, the maximum values of $L$ and $\varepsilon$ decrease with $s$ while the region where they become zero increases.
Another way to investigate the transition from entangled to separable configurations of a given state $\rho$ is to use the concept of robustness [8], which quantifies entanglement with respect to the maximally mixed state $\rho_{\star}$ and may be expressed in terms of the concurrence [14]. Referring to the optimized states of two qubits derived in the previous sections, the main point is to consider the density operator
$\widetilde{\rho}_{X}=(1-\omega)\rho_{X}+\omega\rho_{\star},\quad w\in[0,1].$ (42)
The robustness of the optimized $\rho_{X}$ corresponds to the minimal value of $\omega$, written $\omega_{0}$, for which $\widetilde{\rho}_{X}$ becomes disentangled [8, 14]. The most important feature of $\widetilde{\rho}_{X}$ is that it is also X–shaped. Therefore, the $L$–measure (34) is automatically applicable. A simple calculation yields
$L(\widetilde{\rho}_{B})=2\max\left\{0,(1-\omega)L_{1}(\rho_{B})-\frac{\omega}{4},(1-\omega)L_{2}(\rho_{B})-\frac{\omega}{4}\right\},$ (43)
$L(\widetilde{\rho}_{W_{k}})=2\max\left\{0,(1-\omega)L_{1}(\rho_{W_{k}})-\frac{\omega}{4}\right\},$ (44)
$L(\widetilde{\rho}_{S})=2\max\left\{0,(1-\omega)L_{1}(\rho_{S})-\frac{\omega}{4},(1-\omega)L_{2}(\rho_{S})-\frac{\omega}{4}\right\},$ (45)
where $L_{j}(\rho_{X})$ stands for the first ($j=1$) and second ($j=2$) subtractions occurring in the $L$-measures (35)-(37). Assuming that subtraction $L_{j}(\rho_{X})>0$ defines the value of $L(\rho_{X})$, we obtain the robustness of the optimized states of two qubits:
$\omega_{0}=\frac{L_{j}(\rho_{X})}{L_{j}(\rho_{X})+1/4}.$ (46)
For the Werner–states shown in Figure 8(a), this expression takes the maximal value 4/5 at $q=1$, and minimal value 0 if $q\leq 1/3$. These values correspond to the maximally entangled expression of $\rho_{W_{k}}$ and a completely disentangled form of $\rho_{W_{k}}$, respectively. Similar conclusions are obtained for the two-Bell mixed states shown in Figure 8(b). In turn, the robustness of the generalized Werner–states shown in Figures 9(a) and 9(b) take the maximal values 3/4 and 2/3, respectively.
3.4 Time-dependent systems
@@
Figure 10: (a) The geometric measure $L$ applied to the time-dependent diatomic state $\rho_{\operatorname{2at}}$ for $\gamma=1$ and $n=10$. As time passes, $L$ decreases from 1 to zero and then increases until it is almost equal to 1. This process is repeated over and over again (b) Comparison of $L(\rho_{\operatorname{2at}})$ with the difference $y-x_{0}$. Clearly, inequality $y>x_{0}$ is violated in the same time intervals where $L=0$.
@@
The envelope formed by the minima of S establishes an upper limit for the entanglement measure. To our knowledge, this surprising property of entropy shared by reduced states has not been reported in the literature on the matter.
@@
Figure 11: The envelope formed by the minima of the one-qubit entropy (blue curve) defines an upper bound for the entanglement measures of the time-dependent diatomic state $\rho_{\operatorname{2at}}$ (a) The geometric measure $L$ is red-shaded (b) The entanglement of formation $\epsilon$ is purple-shaded.
@@ 4 Conclusions
With the above results we have constructed a convex representation that facilitates the study of entanglement in two-qubit systems. The description is purely geometric and is based on two of the main characteristics of the X–states:
@@
The $L$-measure reproduces the results of the Hill–Wootters generalized concurrence $C$ for states $\rho_{X}$. That is, our results give $C$ a geometric meaning that is not obvious without the convex structure introduced in this work. What is even better is that, unlike $C$, the geometric measure $L$ distinguishes the rank of states with maximum entanglement.
@@
Note that the universality of two-qubit X–states ensures the utility of our geometric representation for studying entanglement in any configuration of two-qubit states. Indeed, it has been shown that ‘for every two-qubit state there is a two-qubit X–state of same spectrum and entanglement, as measured by concurrence, negativity or relative entropy of entanglement’ [38]. In this sense, no entanglement is lost by avoiding certain coherences of the initial state $\rho$ during the optimization process that we have developed throughout this work. Therefore, a point of the convex set (27), rewritten here as
$\mathcal{S}=\left\{(x,y)\in\mathbb{R}^{2};0\leq|r_{14}|+|r_{23}|\leq\tfrac{1}{2}\right\},$
determines the matrix elements of a density operator $\rho_{X}$ that shares with a two-qubit state (or a family of two-qubit states) $\rho$ both the spectrum and the entanglement. Geometrically, $\mathcal{S}$ is nothing more than the right-triangle shown in Figure 1. Within the geometric formulation presented in this work, the populations of $\rho_{X}$ are assumed to be given. Therefore, the coherences $r_{14}$, $r_{23}$ and their complex conjugates determine not only the eligibility of $\rho_{X}$ and $\rho$ as quantum states but also the possibility of finding entanglement.
The applicability of our method goes far beyond the time-independent systems typically discussed in the literature. We have also studied what happens to entanglement when the system depends explicitly on time. Within the geometric representation, the time-evolution of $\rho_{X}$ describes paths that transit between areas with different entanglement strength, even invading regions where the state becomes separable to return to the entanglement zones, and so on. The latter opens the possibility of manipulating and controlling entanglement in two-qubit systems by solving the quantum inverse problem [54], where one seeks to manipulate systems in order to force them to behave in a particular way. Recent results concerning global and genuine entanglement in three-qubit systems can be found in [18], for other applications see [55, 56, 57, 58, 59, 60, 61].
We have found another outstanding result for the time-dependent system studied here. The envelope formed by the minima of $S$, the entropy of the reduced states, establishes an upper bound for the entanglement measures. To our knowledge, this surprising property has not been reported in the literature on the matter.